# Elastohydrodynamics of 3D chemically active filaments

### v1

@@ 2.1 Local slender phoretic theory
2.2 Elastohydrodynamics
The equations of three-dimensional elastohydrodynamics are solved numerically as described in detail by Walker et al. (2020). In summary, we are considering the Kirchhoff equations, which are the pointwise force and moment balance equations,
$\frac{\partial\mathbf{F}}{\partial s}+\bm{K}=\bm{0},$ (9)
$\frac{\partial\mathbf{M}}{\partial s}$ $+\hat{\bm{t}}\wedge\mathbf{F}+\bm{N}=\bm{0}.$ (10)
where $\mathbf{F}$ is the internal force acting on the cross-section of the rod at $s$, and $\mathbf{M}$ is the bending moment. We identify the force and torque densities, $\bm{K}$ and $\bm{N}$, as arising from the viscous tractions associated with the instantaneous phoretic slip flow, which is computed using the local slender phoretic theory described in section 2.1. This is then coupled to the elastohydrodynamic problem analogously to how a background flow is accounted for by Walker et al. (2020).
@@ 2.3 Numerical implementation
where the vector $\bm{\Theta}$ records the locally defined Euler angles and the position of one end of the slender body, from which the current configuration can be uniquely determined. The linear operators $B$, $A$, and $Q$ encode various aspects of the theory: $Q$ converts from the angular parameterisation to the lab-frame coordinates of the endpoints of the $N$ discrete segments; $A$ uses the resistive force theory of Hancock (1953); Gray & Hancock (1955) and the resistive torque theory of Walker et al. (2023) to relate linear and angular velocities to the forces and torques on the slender body; $B$ encodes the governing equations of force and torque balance on the slender body. Accordingly, the vector $\bm{R}$ records any applied forces and torques on the body. In this case, $\bm{R}$ includes elastic restoring moments and the forces and torques caused by the phoretic slip velocity.
@@
which is defined in terms of the fluid viscosity $\mu$, the half-length $l$ of the slender body, the bending stiffness $EI$ of the body, and a characteristic timescale $T$. Here, we take $T=Dl^{2}/AMr_{f}$ as the characteristic timescale set by Slender Phoretic Theory. The elastohydrodynamic number increases with chemical activity and mobility, and decreases with increasing filament bending stiffness, so that high $E_{h}$ corresponds to strong slip flows or very flexible filaments.
The linear system defined by $BAQ$ is square, so that one may readily solve for $\dot{\bm{\Theta}}$ to yield evolution equations for the position and configuration of the slender body. These are implemented in MATLAB and solved using the adaptive, implicit timestepping scheme ode15s (Shampine & Reichelt, 1997). In adapting the implementation of Walker et al. (2020), we are implicitly assuming that the slender body is inextensible, unshearable and moving in a Newtonian fluid at vanishing Reynolds number. We also employ the Euler–Bernoulli constitutive relation of the Kirchhoff formalism with a Poisson ratio of unity (Antman, 2005; Nizette & Goriely, 1999).
@@ 3 Results
In this section, we present an anthology of results for the intricate dynamics of chemoelastohydrodynamic filaments, through long-time, three-dimensional simulations for different chemical activity profiles. We focus on activity profiles that decay to zero at the filament tips, to avoid unphysical large loads that might arise from the breakdown of the slenderness assumptions of SPT near the ends, $s=\pm 1$. We thus consider two key axisymmetric activity profiles,
[…]
which represent (a) a symmetric smooth activity profile, and (b) an antisymmetric smooth activity profile. In the rigid analogue, these profiles would behave as (a) Saturn and (b) Janus rods, respectively. Unless otherwise stated, the initial configuration is an S-shaped planar curve, which is deformed from the elastic rest state of a perfectly straight filament. The activity and tangential slip velocity for these filaments in the straight rod configuration is shown in Figure 2 for a mobility $\mathcal{M}=+1$. Note that, when straight, the symmetric activity (a) generates a compressive flow along the whole filament, while the antisymmetric example (b) has regions of both compression and extension.
@@
Figure 2: The chemoelastic filaments considered. All shapes are prolate spheroidal, with cross-sectional radius $\rho=\sqrt{1-s^{2}}$. The considered activities are $\mathcal{A}=\sqrt{1-s^{2}}$ and $\mathcal{A}=\sin(\pi s)$. (a) Representations of the activity and shape of the filaments in their elastic rest state. (b) The prescribed activity and (c) the slip velocity calculated from local SPT for the square root (blue) and sinusoidal (green) activities, when $\mathcal{M}=+1$. The symmetrically-patterned filament experiences a compressive flow, while the antisymmetric filament has regions of compression and extension.
We focus on the effect that varying the elastohydrodynamic number $E_{h}$ has on the swimming dynamics, in a similar manner to studies based upon follower forces (e.g. Laskar & Adhikari, 2017; De Canio et al., 2017; Fily et al., 2020; Man & Kanso, 2019; Clarke et al., 2024; Schnitzer, 2025; Altunkeyik et al., 2025) and force distributions (e.g. Laskar et al., 2013; Manna et al., 2017; Stein et al., 2021; Lough et al., 2023). We vary $E_{h}$ within the range $E_{h}=1000$ (very stiff, low activity) to $E_{h}=100,000$ (very deformable, high activity), looking for characteristic behaviours. In particular, we are interested in examining the transitions between different dynamic regimes; for instance, the initial buckling of an elastic filament, and transitions between planar, periodic and chaotic behaviours that may arise as the filament becomes more deformable (or, equivalently, as the active forces increase).
@@ 3.1 Saturn-like activity: Fore-aft symmetric profiles
A perfectly rigid, straight rod with an activity $\mathcal{A}=\sqrt{1-s^{2}}$ does not move and generates a straining flow where fluid is pulled inwards from both poles of the rod. We therefore expect the filament to be under a compressive loading, and thus exhibit buckling for sufficiently high values of the elastohydrodynamic number $E_{h}$. Similar phenomena are observed for filaments driven by non-conservative ‘follower’ forces, as discussed in the introduction.
@@
We consider the dynamics of these filaments, beginning at stiff filaments with low values of $E_{h}\approx 1000$, and increasing to $E_{h}=100,000$. This increase in $E_{h}$ can be considered as increasing the deformability (decreasing stiffness) with fixed activity, or equivalently increasing the activity with fixed stiffness.
@@ 3.1.1 Planar buckling: low $E_{h}$
Once $E_{h}$ is above a critical value, close to $E_{h}\approx 2000$, we observe a transition in behaviour: the filament deforms to form a clear U-shape, and moves forwards on a ballistic, planar trajectory away from the tips of the U-shape. An example trajectory is shown in Figure 3a for $E_{h}=5000$. This transition is reminiscent of Euler buckling; starting from a small perturbation from a straight rod, the compressive forcing due to the chemical activity is large enough to buckle the filament into its first Euler mode, and the resulting shape asymmetry generates concentration gradients between the inside and outside of the curve that drive swimming. This buckling instability unlocks new modes of swimming via shape change from I-shaped pumps to U-shaped swimmers (Sharan et al., 2021; Montenegro-Johnson, 2018).
In this U-shaped swimming, we soon observe a settling towards a steady propelling state, which varies with $E_{h}$. For example, as $E_{h}$ increases, so does the curvature of the filament at steady state, as shown in Figure 3b. This can be explained, for example, by considering the same compressive load being applied to a relatively more flexible filament, promoting further bending. With both ends of the filament pointing more directly backwards, more propulsive force is aligned with the direction of motion; the filament must remain force free and so it moves more quickly to balance this force via viscous drag, which can be seen in Figure 3c. Long-time simulations indicate that that these planar trajectories are likely stable for elastohydrodynamic numbers as large as $E_{h}\approx 14,000$.
@@
Figure 3: Planar ballistic motion of a filament with symmetric activity $\mathcal{A}=\sqrt{1-s^{2}}$ at low $E_{h}$. (a) An example trajectory with $E_{h}=5000$ up to $t=40$, with time progressing from dark to light. (b) The maximum curvature at steady state as a function of $E_{h}$, with insets showing example filament shapes. (c) The ballistic speed increases with $E_{h}$ before plateauing.
@@ 3.1.2 Onset of three-dimensional motion: moderate $E_{h}$
As the elastohydrodynamic number increases further, we see an interesting transition from the previous planar motion to fully three-dimensional dynamics, out-of-the-plane. Beginning with our initial planar S-shape perturbation, we observe this clearly within time $t=100$ from approximately $E_{h}=14,400$. Additional simulations, with an initial fully three-dimensional (non-planar) helical initial condition suggest that this transition from steady U-shaped planar translation to three-dimensional motion may occur as low as $E_{h}=14,000$.
In simulations just above this threshold, the filament initially travels in a planar U-shaped translating mode for a period of time, but then small amounts of out-of plane motion cause it to begin revolving about its direction of motion as it progresses forwards. As $E_{h}$ increases further, the out-of-plane spinning increases in angular frequency, the filament begins to wobble, and for higher values from $E_{h}\approx 14,600$, we see a sharp, non-planar turn, followed by coherent ballistic helical motion out of the plane. The ballistic helical motion comprises a translational component, and a spin about the axis of translation, with a small wobble. Long run-time simulations indicate that the resulting ‘wobbly helix’ trajectory is stable. This behaviour is shown for $E_{h}=15,000$ in Figure 4.
@@
Figure 4: Above a threshold $E_{h}\approx 14,000$, trajectories deviate out-of-plane. Here, an example is shown for $E_{h}=15,000$. (a) Trajectory snapshots up to time $t=100$, showing the initially planar trajectory begin to spin, and then sharply turn to settle into a near-helical trajectory. Heatmaps of (b) the filament curvature and (c) torsion demonstrate a clear transition in geometry and dynamics.
@@ 3.1.3 Stable periodic orbits: high $E_{h}$
The above-described out-of-plane helical trajectory undergoes a critical transition to a trapped periodic state as $E_{h}$ increases further. For example, after an initial transient, the filament follows a tight helical path for $E_{h}=16,000$, but beyond $E_{h}=17,000$, the early-time planar translation deviates quickly to a tight, circular path that does not translate. An example of this is shown in Figure 5 for $E_{h}=23,000$. We refer to this state as ‘pinwheeling’.
This pinwheeling behaviour, once exhibited, seems to be a stable attractor. As $E_{h}$ increases, the early-time dynamics prior to pinwheeling goes through a range of ballistic trajectories, including planar translation, flapping back and forth and more chaotic-like motion, but these trajectories soon transition to the stable orbiting motion. These orbits appear to exist up to $E_{h}\approx 33,000$.
During pinwheeling, the shape of the filament appears to be fixed, as shown in Figure 5c. It is symmetric about the midpoint, and takes the form of a U-shape that is bent out-of-plane near the tips. This bend becomes more pronounced as $E_{h}$ increases, resulting in a tighter pinwheeling trajectory, i.e. motion on a small circular attractor.
@@
Figure 5: Stable periodic orbits exist at intermediate $E_{h}$. (a) Snapshots of a stable ‘pinwheeling’ trajectory, showing its characteristic tight circular motion for $E_{h}=23,000$. (b) A heatmap of the curvature shows an initial period of symmetric periodic flapping motion that gives way to the stable fixed shape of pinwheeling. (c) While orbiting, the filament is in an approximately fixed configuration, that can be described as a folded or bent over U-shape. As $E_{h}$ increases, the U-shape becomes increasingly non-planar, resulting in tighter circular motion.

### v2

@@ 2.2 Local slender phoretic theory
2.3 Resistive force theory
The phoretic slip flow induces bulk flow in the surrounding fluid, acting as a boundary condition to the Stokes equations for a viscous fluid,
$\mu\nabla^{2}\bm{u}=\nabla p,\qquad\nabla\cdot\bm{u}=0,$
where $\mathbf{u}(\mathbf{x},t)$ is the fluid velocity, $p(\mathbf{x},t)$ is the fluid pressure and $\mu$ is the fluid’s dynamic viscosity. The flow is assumed to decay to $\mathbf{u}=\bm{0}$ in the far field, and satisfies a slip boundary condition on the surface of the body.
To model the Stokes flow around the slender filament, we use resistive force theory (Hancock, 1953; Gray & Hancock, 1955). This is a local slender body theory for viscous flow that decomposes the local motion of the filament (relative to a background or slip flow) into components that are locally tangential and normal to the centreline, linearly relating the leading order force density and velocity of each via different drag coefficients. Here, the appropriate velocity is the surface velocity in the lab frame (the frame in which $\bm{u}\to 0$ far from the filament), which is $\bm{u}_{motion}+\bar{\bm{v}}_{slip}$, where $\bm{u}_{motion}$ corresponds to the local velocity of the slender body. Note that this captures the notion that slip velocities act in the opposite direction to the motion that they drive, so that a rigid body moves in the opposite direction to the mean slip velocity on its surface (in the absence of other effects). We decompose this combined quantity into $\bm{u}_{\parallel}$ and $\bm{u}_{\perp}$, the components in the tangential and normal directions, respectively. The leading order force density is linearly related to this local relative velocity via
$\bm{f}_{\parallel}=-C_{\parallel}\bm{u}_{\parallel},\qquad\bm{f}_{\perp}=-C_{\perp}\bm{u}_{\perp},$ (9)
where $\bm{f}_{\parallel},\bm{f}_{\perp}$ denote the force density applied on the filament in directions tangential to the centreline and normal to it. The drag coefficients are
$C_{\parallel}=\frac{2\pi\mu}{\log(4/\rho\epsilon)-0.5},\qquad C_{\perp}=\frac{4\pi\mu}{\log(4/\rho\epsilon)-0.5}.$ (10)
The resistive torque theory of Walker et al. (2023) links local angular velocity (such as that about the local tangent) to local torque via an additional resistive coefficient. Following this, here we take the local applied torque per unit length to be $4\pi\epsilon^{2}\rho(s)^{2}\mu\bm{\Omega}$ for angular velocity $\bm{\Omega}$ of the slender body. This application of resistive force theory is valid for the slip flows generated by chemically-propelled filaments, since local slender body theories are consistent with boundary conditions that enforce rigid-body motion locally, as is the case here (Walker et al., 2023).
2.4 Kirchoff rod equations
We model the deformation of the filament using the Kirchhoff equations, which are the pointwise force and moment balance equations,
$\frac{\partial\mathbf{F}}{\partial s}+\bm{f}=\bm{0},$ (11)
$\frac{\partial\mathbf{M}}{\partial s}$ $+\hat{\bm{t}}\wedge\mathbf{F}+\bm{\tau}=\bm{0}.$ (12)
where $\mathbf{F}$ is the internal force acting on the cross-section of the rod at $s$, and $\mathbf{M}$ is the bending moment. We identify $\bm{f}$ and $\bm{\tau}$ as the force and torque densities due to the viscous tractions generated by the flow; these are computed using the resistive force framework of section 2.3.
We impose conditions of zero force and torque at the filament ends, so that
$\mathbf{F}(-1)=\mathbf{F}(1)=\mathbf{M}(-1)=\mathbf{M}(1)=0.$ (13)
Integrating eqs. 11 and 12 leads to the governing equations
$\int_{-1}^{1}$ $\bm{f}(\tilde{s})~\mathrm{d}\tilde{s}=0,$ (14)
$\int_{s}^{1}[\bm{X}(\tilde{s})-\bm{X}(s)]$ $\times\bm{f}(\tilde{s})+\bm{\tau}(\tilde{s})~\mathrm{d}\tilde{s}=\mathbf{M}(s).$ (15)
The bending moment is given by the constitutive equation
$\mathbf{M}=EI\left(\kappa_{1}\bm{d}_{1}+\kappa_{2}\bm{d}_{2}+\frac{\kappa_{3}}{1+\nu}\bm{d}_{3}\right),$ (16)
where $\bm{d}_{i}$ are a right-handed orthonormal director basis with $\bm{d}_{3}=\hat{\bm{t}}$, and $\kappa_{i}$ are the components of the twist vector such that $\partial\bm{d}_{i}/\partial s=\bm{\kappa}\times\bm{d}_{i}$. The material parameters are captured by the bending stiffness, $EI$, and the Poisson ratio, $\nu$; for simplicity, we assume both of these are constant throughout the filament, with $\nu=1$ (Antman, 2005; Nizette & Goriely, 1999).
@@ 2.5 Numerical implementation
where the vector $\bm{\Theta}$ records the locally defined Euler angles and the position of one end of the slender body, from which the current configuration can be uniquely determined. The linear operators $B$, AA, and $Q$ encode various aspects of the theory: $Q$ converts from the angular parameterisation to the lab-frame coordinates of the endpoints of the $N$ discrete segments; $A$ uses the resistive force theory of Hancock (1953); Gray & Hancock (1955) and the resistive torque theory of Walker et al. (2023) to relate linear and angular velocities to the forces and torques on the slender body; $B$ encodes the governing equations of force and torque balance on the slender body. Accordingly, the vector $\bm{R}$ records any applied forces and torques on the body. In this case, $\bm{R}$ includes elastic restoring moments and the forces and torques caused by the phoretic slip velocity.
@@
which is defined in terms of the fluid viscosity $\mu$, the half-length $l$ of the slender body, the bending stiffness $EI$ of the body, and a characteristic timescale $T$. Here, we take $T=Dl^{2}/AMr_{f}$ as the characteristic timescale set by slender phoretic theory. The elastohydrodynamic number therefore represents the ratio of the elasto-viscous timescale, $t_{ev}\propto\mu l^{4}/EI$, to the flow timescale $T$. It increases with chemical activity and mobility, and decreases with increasing filament bending stiffness, so that high $E_{h}$ corresponds to strong slip flows or very flexible filaments.
The linear system defined by $BAQ$ is square, so that one may readily solve for $\dot{\bm{\Theta}}$ to yield evolution equations for the position and configuration of the slender body. These are implemented in MATLAB and solved using the adaptive, implicit timestepping scheme ode15s (Shampine & Reichelt, 1997). In adapting the implementation of Walker et al. (2020), we are implicitly assuming that the slender body is inextensible, unshearable and moving in a Newtonian fluid at vanishing Reynolds number..
@@ 3 Results
In this section, we present an anthology of results for the intricate dynamics of chemoelastohydrodynamic filaments, exploring long-time, three-dimensional simulations for different chemical activity profiles. We focus on a prolate spheroidal geometry, $\rho(s)=\sqrt{1-s^{2}}$, and activity profiles that decay to zero at the filament tips. Our theory is valid for more general profiles and chemical patterning than this (see e.g.  Katsamba et al., 2020); however, we focus on these examples to avoid more technical hurdles that might arise in choosing an appropriate discretisation to account for boundary layer effects (Katsamba et al., 2022) or from the breakdown of the slenderness assumptions of SPT near the ends, $s=\pm 1$, for shapes that taper too rapidly. We thus consider two key axisymmetric activity profiles,
$\text{(a) }~\mathcal{A}=\sqrt{1-s^{2}},\hskip 28.45274pt\text{(b) }~\mathcal{A}=\sin{\pi s},$ (19)
which represent (a) a symmetric smooth activity profile, and (b) an antisymmetric smooth activity profile. In the rigid analogue, these profiles would behave as (a) Saturn and (b) Janus rods, respectively. Unless otherwise stated, the initial configuration is an S-shaped planar curve, which is deformed from the elastic rest state of a perfectly straight filament. The activity and tangential slip velocity for these filaments in the straight rod configuration is shown in Figure 2 for a mobility $\mathcal{M}=-1$. Note that, when straight, the symmetric activity (a) generates an extensional flow in the surrounding fluid, from the centre to the tips; this causes compression in the elastic filament as the phoretic surface is propelled against the slip flow. For the antisymmetric example (b), there are regions of both compression and extension.
@@
Figure 2: The chemoelastic filaments considered. All shapes are prolate spheroidal, with cross-sectional radius $\rho=\sqrt{1-s^{2}}$. The considered activities are $\mathcal{A}=\sqrt{1-s^{2}}$ and $\mathcal{A}=\sin(\pi s)$. (a) Representations of the activity and shape of the filaments in their elastic rest state. (b) The prescribed activity and (c) the slip velocity calculated from local SPT for the square root (blue) and sinusoidal (green) activities, when $\mathcal{M}=-1$. The symmetrically-patterned filament generates a surrounding extensional flow that applies compressive forces on the filament, while the antisymmetric filament has regions of both compression and extension.
We focus on exploring the effect that varying the elastohydrodynamic number $E_{h}$ has on the swimming dynamics, in a similar manner to studies based upon follower forces (e.g. Laskar & Adhikari, 2017; De Canio et al., 2017; Fily et al., 2020; Man & Kanso, 2019; Clarke et al., 2024; Schnitzer, 2025; Altunkeyik et al., 2025) and force distributions (e.g. Laskar et al., 2013; Manna et al., 2017; Stein et al., 2021; Lough et al., 2023). We vary $E_{h}$ within the range $E_{h}=1000$ (very stiff, low activity) to $E_{h}=100,000$ (very deformable, high activity), looking for characteristic behaviours. In particular, we are interested in examining the transitions between different dynamic regimes; for instance, the initial buckling of an elastic filament, and transitions between planar, periodic and chaotic behaviours that may arise as the filament becomes more deformable (or, equivalently, as the active forces increase).
@@ 3.1 Saturn-like activity: Fore-aft symmetric profiles
A perfectly rigid, straight rod with an activity $\mathcal{A}=\sqrt{1-s^{2}}$ does not move and generates a straining flow where fluid is pushed outwards towards the poles of the rod. We therefore expect the filament to be under a compressive loading and, thus, exhibit buckling for sufficiently high values of the elastohydrodynamic number $E_{h}$. Similar phenomena are observed for filaments driven by non-conservative follower forces, as discussed in the introduction.
@@
We consider the dynamics of these filaments, beginning with stiff filaments that have low values of $E_{h}\approx 1000$, and increasing to $E_{h}=100,000$. This increase in $E_{h}$ can be considered as increasing the deformability (decreasing stiffness) with fixed activity, or equivalently increasing the activity with fixed stiffness.
@@ 3.1.1 Planar buckling: low $E_{h}$
Once $E_{h}$ is above a critical value, close to $E_{h}\approx 2000$, we observe a transition in behaviour: the filament deforms to a clear U-shape, and moves forwards on a ballistic, planar trajectory away from the tips of the U-shape. An example trajectory is shown in Figure 3a for $E_{h}=5000$. This transition is reminiscent of Euler buckling; starting from a small perturbation from a straight rod, the compressive forcing due to the chemical activity is large enough to buckle the filament into its first Euler mode, and the resulting shape asymmetry generates concentration gradients between the inside and outside of the curve that drive swimming. This buckling instability unlocks new modes of swimming via shape change from I-shaped pumps to U-shaped swimmers (Sharan et al., 2021; Montenegro-Johnson, 2018).
In this U-shaped swimming, we observe a settling towards a steady propelling state, which varies with $E_{h}$. For example, as $E_{h}$ increases, so does the curvature of the filament at steady state, as shown in Figure 3b. This can be explained, for example, by considering the same compressive load being applied to a relatively more flexible filament, promoting further bending. With both ends of the filament pointing more directly backwards, more propulsive force is aligned with the direction of motion; the filament must remain force-free and so it moves more quickly to balance this force via viscous drag, which can be seen in Figure 3c. Long-time simulations indicate that these planar trajectories are likely stable for elastohydrodynamic numbers as large as $E_{h}\approx 21,000$. Even from out-of-plane helical initial conditions, planar behaviour seems to be a stable attractor for $E_{h}$ up to approximately $20,000$.
@@
Figure 3: Planar ballistic motion of a filament with symmetric activity $\mathcal{A}=\sqrt{1-s^{2}}$ at low $E_{h}$. (a) An example trajectory with $E_{h}=5000$ up to $t=40$, with time progressing from dark to light. (b) The maximum curvature at steady state as a function of $E_{h}$, with insets showing example filament shapes, computed until $t=10$. (c) The ballistic speed increases with $E_{h}$ before plateauing.
@@ 3.1.2 Onset of three-dimensional motion: moderate $E_{h}$
As the elastohydrodynamic number increases further, we observe a transition from planar motion to fully three-dimensional dynamics, with the filament leaving the initial plane of motion. Beginning with our initial planar S-shape perturbation, we observe this clearly within time $t=100$ from $E_{h}\approx 21,500$.
Just above this threshold, the initially planar filament adopts a U-shaped translating mode for a period of time, but then small amounts of out-of plane motion cause it to begin revolving about its direction of motion as it progresses forwards. As $E_{h}$ increases further, above approximately $23,000$, the out-of-plane spinning increases in angular frequency, the filament begins to wobble and, for higher values of $E_{h}$, we observe a sharp, non-planar turn followed by coherent ballistic helical motion out of the plane. This latter mode comprises a translational component and a spin about the axis of translation, with a small wobble. Extended simulations suggest that the resulting ‘wobbly helix’ trajectory is stable when it exists. This behaviour is shown for $E_{h}=23,000$ in Figure 4.
@@
Figure 4: At long times and above a threshold of $E_{h}\approx 21,500$, trajectories deviate out-of-plane. Here, an example is shown for $E_{h}=23,000$. (a) Trajectory snapshots up to time $t=100$, with time progressing from dark to light, showing the initially planar trajectory begin to spin, and then sharply turn to settle into a near-helical trajectory. Heatmaps of (b) the filament curvature and (c) torsion demonstrate a clear transition in geometry and dynamics.
@@ 3.1.3 Stable periodic orbits: high $E_{h}$
The out-of-plane helical trajectories described above undergo a critical transition to an apparently periodic state as $E_{h}$ increases slightly further. For example, after initial transients, the tight helical path of $E_{h}=23,000$ is lost by $E_{h}=24,000$, whereafter the early-time planar translation deviates quickly to a tight, circular path that does not translate over long timescales. An example of this is shown in Figure 5 for $E_{h}=32,000$. We refer to this state as ‘pinwheeling’.
This pinwheeling behaviour, once exhibited, seems to be a stable attractor. As $E_{h}$ increases, the early-time dynamics prior to pinwheeling goes through a range of ballistic trajectories, including planar translation, flapping back and forth and more chaotic-like motion, but these trajectories soon transition to the stable orbiting motion. These orbits appear to exist up to $E_{h}\approx 48,000$.
During pinwheeling, the shape of the filament appears to be approximately fixed; these shapes are illustrated in Figure 5c. The filament is symmetric about the midpoint and takes the form of a U-shape that is bent out-of-plane near the tips. This bend becomes more pronounced as $E_{h}$ increases, resulting in a tighter pinwheeling trajectory, i.e. motion on a small circular attractor.
@@
Figure 5: Stable periodic orbits exist at intermediate $E_{h}$. (a) Snapshots of a stable ‘pinwheeling’ trajectory, with time progressing from dark to light, showing its characteristic tight circular motion for $E_{h}=32,000$. (b) A heatmap of the curvature shows an initial period of symmetric periodic flapping motion that gives way to the stable fixed shape of pinwheeling. (c) While orbiting, the filament is in an approximately fixed configuration, that can be described as a folded or bent over U-shape. As $E_{h}$ increases, the U-shape becomes increasingly non-planar, resulting in tighter circular motion.